# Deep learning framework for action prediction reveals multi-timescale locomotor control

### v2

@@ Abstract
Modeling movement in real-world tasks is a fundamental scientific goal for motor control, biomechanics, and rehabilitation engineering. However, existing models and their simplifying assumptions such as linear and fixed timescale mappings do not generalize to real-world contexts. Here, we develop a deep learning-based framework for action prediction with architecture-dependent trial embedding, outperforming traditional models across multiple contexts (walking and running, treadmill and overground, varying terrains) and input modalities (multiple body states, gaze). We find that neural network architectures with flexible input history-dependence like GRU and Transformer perform best overall. By quantifying the model’s predictions relative to an autoregressive baseline, we identify context- and modality-dependent timescales. There is greater reliance on fast-timescale predictions in complex terrain, gaze predictions precede body state predictions, and full-body state predictions precede center-of-mass-relevant predictions. This deep learning framework for action prediction provides quantifiable insights into the control of complex movements and can be extended to other actions, contexts, and populations.
1 Introduction
@@
Modeling human movement in real-world settings remains a major scientific challenge for motor control, biomechanics, and rehabilitation engineering. Data-driven models of human movement can inform predictive simulations [1, 2], characterize neuromotor disorders [3, 4], and can be incorporated into wearable robots to anticipate the user’s actions [5, 6]. Data-driven models widely used to characterize locomotor control rely on simplifying assumptions such as linear, fixed-timescale mappings or are limited to laboratory-constrained tasks. Such models have successfully explained how center-of-mass (CoM) kinematics predict foot placement [7, 8, 9, 10] and ground reaction forces [2, 11] during treadmill locomotion, but they have not been shown to generalize to real world settings. During real world movements, people exhibit more variable behaviors [12, 13], and the control strategies may depend on the nonlinear history of multiple task-relevant inputs [14]. Addressing this requires a framework capable of predicting future actions given time-varying inputs from different modalities, such as body state and visual gaze, across contexts. Artificial neural networks have shown significant promise in predicting foot placement actions, as deep learning models can accurately capture the temporal and nonlinear statistics at play [15]. However, previous neural network models of locomotion test a single architecture [16, 17, 18], use a single input modality [16, 17] or are confined to a single laboratory-constrained setting like treadmill walking [18, 19]. Here, we build nonlinear models comparing multiple deep learning architectures, across real world contexts, and training on different input modalities to predict future foot placements.
@@
Biological motor control is time-varying and context-dependent, requiring adaptive strategies to meet environmental, task-specific, and physiological demands [20]. For example, during locomotion, humans possess a remarkable ability to continuously modulate their actions in diverse “contexts" i.e. environment-specific circumstances such as uneven terrain or obstacle-avoidance [21, 22]. This ability is supported by dynamic sensorimotor control strategies that can be deployed both on a fast timescale such as during each walking step, or on a longer timescale over multiple steps [22, 23]. For instance, walking on uneven terrain requires rapid adjustments to foot placement [24], as well as persistent motor plans involving path or speed selection [25]. In fact, multiple timescales in motor control have been hypothesized to reflect distinct neural computations [26] or distinct functional goals [27], so quantifying multi-timescale control strategies from real world data has theoretical value. Despite this, we lack meaningful approaches to quantify multi-timescale control in real world contexts. Here, we put forth the relative predictive power of nonlinear action prediction models relative to a baseline as a generalizable way to quantify control timescales, discovering a context-dependent tradeoff between fast- and slow-timescale predictions.
@@
Biological sensorimotor control is achieved by dynamically mapping multi-sensory inputs to motor commands, enabling adaptive control in changing contexts [28, 29]. During locomotion, the control of foot placement is inherently multimodal, and prior work has examined the role of individual input modalities such as vision [24] or low-dimensional body states on control [7]. Visual input is crucial for anticipatory locomotor control when navigating obstacles or selecting stable footholds on complex terrains [13, 30], while joint-level or CoM-based feedback informs muscle responses to transient perturbations [31, 32, 2]. Vestibular input is also thought to guide foot placement, particularly in a phase-dependent manner [33]. Despite the existence of multiple input modalities, existing foot placement control models assume a linear fixed control timescale with low-dimensional body states as an input [7, 34], thereby neglecting the temporal dynamics of how and when each modality contributes to control, and how different modalities compare to one another. To address these gaps, we propose a data-driven framework that systematically evaluates the contributions of multiple input modalities — including CoM-relevant kinematics, full-body kinematics, and gaze fixations — to predicting future foot placement. By comparing the predictive power of each modality to a swing foot dynamics based baseline model, we quantify the timescale and magnitude of modality-dependent foot placement. Our framework reveals at which phase of the gait cycle each input modality become critical for informing future foot placement.
@@
In this study, we develop deep learning-based models to predict future actions from multimodal input time series and test them in diverse locomotor tasks. By comparing different model architectures, we find that architectures with flexible input history-dependence generalize better across real-world contexts and input modalities. We characterize the control timescale of a given nonlinear action prediction model using its predictive power relative to a baseline model, to help reveal how these timescales are shaped by the context and the input modality. Our work has implications for characterizing, modeling, and predicting real world movements and the multi-timescale control strategies hidden in them. These nonlinear action prediction models and timescale analyses can inform the development of human-aware wearable robot control and more human-like physics-based simulations of motor behavior in the real world [35, 36].
2 Materials and Methods
@@
2.1 Data Description and Processing
@@ 2.2 Modeling Framework
Figure 1: Overview of the data-driven modeling framework. A Inputs contain hypothesized modalities such as CoM-relevant kinematics, swing-foot kinematics, full-body kinematics, and gaze fixations (defined by the intersection of the gaze vector with the ground plane). The output is foot placement. B Foot placement output is defined by the lateral and fore-aft foot placement relative to the opposite foot at its previous heel strike. C The input is a time series spanning from three strides prior to touchdown to a specific gait phase within the previous stride. The output corresponds to foot placement at heel strike. We denote the ending gait phase of the time series relative to one stride prior to touchdown as $\phi$, discretized and sampled from $n=21$ equally spaced phases within $0\leq\phi\leq 1$. D The network architecture of the nonlinear models. The input consists of a time series, an L/R flag, and a trial ID. The time series and L/R flag are concatenated, while the trial ID is embedded to provide additional trial-specific information. This embedding influences the model prediction differently depending on the nonlinear model architecture, as detailed in section 2.2.1.
2.2.1 Nonlinear models
@@ 2.2.2 Linear models
2.3 Model Evaluation and Baseline Model
@@
Nested cross-validation was employed to evaluate the models’ performance and hyperparameter tuning, ensuring an unbiased estimate of model performance and preventing data leakage [47]. The dataset was initially split into five equal folds. In each iteration of the outer loop, 80% of the data was allocated for training and validation, while the remaining 20% was held out as the testing set. The training/validation set was further subdivided into five inner folds. Four inner folds were used to train the model, and the remaining fold was used for early stopping, with a patience of 50 epochs. This process was repeated for every combination of hyperparameters, with parameter updates for nonlinear models performed using the ADAM optimizer [48] and a learning rate of 0.001. We used mean squared error (MSE) as the loss function. After each training session, the model’s performance was evaluated on the outer testing fold. The outer loop was repeated five times, ensuring that each data point in the original dataset was used once for testing and four times for training/validation.
@@
Optimal hyperparameters for each modality and dataset were first identified using grid search (or random search for large hyperparameter spaces exceeding 100 configurations; see Appendix Hyperparameter Tuning for details). These hyperparameters were then used in the nested cross-validation process. Aggregated predictions from all outer folds, covering 100% of the data points, were used to compute the explained variance ($R^{2}$) of foot placement predicted by the input modality. Under this nested cross-validation framework, it is possible, though unlikely, to obtain negative $R^{2}$ values, which occurs when the model fails to generalize to the testing dataset, performing worse than simply predicting the average foot placements (which would result in $R^{2}=0$). In such cases, the model’s predictive power is minimal or nonexistent. To prevent overfitting, we used early stopping and dropout during training for nonlinear models, whereas $L2$ regularization is used to prevent overfitting to historical data or high-dimensional input space for linear models. Since neural networks are more complex than linear models, they inherently exhibit lower bias and higher variance. To mitigate this higher variance of neural networks and to better interpret model behavior across gait phases, we smooth the $R^{2}$ curves using local regression (LOWESS) and cubic spline.
@@
To evaluate model performance, we assigned a score to each model based on the Root Mean Squared Error (RMSE). As the time window increases, the input contains accumulative information about the foot placement thereby enhancing the models’ predictive power. We used the RMSE curve achieved by the swing-foot kinematics as a baseline, reflecting the predictive information inherently contained in the swing foot dynamics [49]. We define the model score as the RMSE as defined in Section 2.4). The normalized model score is then computed relative to the model with the highest score. Formally, let $\text{RMSE}(m,\phi)$ denote the RMSE of model $m$ at gait phase $\phi$ and let $c$ denote the critical gait phase. We define the model score $s_{m}$:
@@
2.4 Quantifying the prediction timescale
The predictive power of foot placement using different input modalities across the gait cycle provides insight into the timescale of the control strategy. For an input modality to be considered a strong predictor, it must outperform the baseline and demonstrate an earlier and larger increase in predictive power. To evaluate this, we identify the gait phase where the input modality achieves its maximum performance over the baseline, measured by the relative RMSE gap (Figure 2A). Beyond this phase, the input modality’s contribution diminishes, as the predictive power of the swing foot itself increases at a faster rate.
@@ 2.4 Quantifying the prediction timescale
Figure 2: Identification of key timepoints for dynamics and prediction. A The relative RMSE between modality-based predictions and the baseline prediction as a function of gait phases. The timing of the peak difference indicates the phase where the input modality is the strongest predictor. B An example velocity curve along the locomotor axis, with the 5% of peak velocity threshold used to determine the onset of the swing phase. C The baseline $R^{2}$ curve, showing a breakpoint where $R^{2}$ increases significantly. The timing of this breakpoint is identified as the phase where the slope exceeds the average slope.
@@
We analyze the baseline $R^{2}$ curve to determine the time at which the swing foot kinematics begins predicting substantially the future foot placement. The $R^{2}$ curve typically shows a non-decreasing trend as the gait cycle progresses, eventually reaching $R^{2}=1$ at the final timestep. A notable “breakpoint” in the $R^{2}$ curve is the phase at which the predictive power of the swing foot kinematics sharply rises. To systematically identify this breakpoint, we calculate the average slope of the $R^{2}$ curve, defined as $r=R^{2}(1)-R^{2}(0)$, and locate the gait phase where the slope reaches its maximal value. We then trace backward from this gait phase to find the first point where the slope falls below the average slope, marking the critical transition phase (Figure 2C). This approach provides a robust method for identifying key transitions in the gait cycle and aligns with the underlying dynamics of gait and motor control.
3 Results
@@
In this section, we present our results on the use of nonlinear action prediction models to understand the influence of the environmental context and input modality on future foot placement. First, we demonstrate that nonlinear models, particularly GRUs, significantly outperform linear models in more challenging contexts. During overground walking, we find that subjects exhibit a tradeoff between fast and slow timescale predictions, with the magnitude of the tradeoff varying across different terrains. By comparing the predictive power of input modalities to a baseline model, we reveal that control timescales are both modality- and context-dependent. Finally, we discover a gait phase mid-swing when the swing foot dynamics becomes useful for foot placement prediction.
@@
3.1 Models with flexible nonlinear input dependence outperform linear models
@@
We evaluated the performance of nine models — LSTM, GRU, FCNN, TCN, Transformer, Linear instance with (LI2) and without (LI) L2 regularization, Linear history with (LH2) and without (LH) L2 regularization — on their ability to predict foot placement across various tasks and input modalities. The tasks included treadmill walking, treadmill running, and overground walking on various terrains, while the input modalities consisted of CoM-relevant kinematics, full-body kinematics, swing-foot kinematics and gaze fixations. The model score was evaluated based on the Root Mean Squared Error (RMSE) relative to the optimal model at each gait phase, and then normalized by the highest-scoring model. Section 2.3 details the evaluation metric.
@@
[…]
Table 1: Normalized Model Scores across Tasks and Input Modalities. Models were evaluated on treadmill walking, treadmill running, and overground walking tasks with varying terrain. The input modalities (M) were: CoM-relevant kinematics (C), full-body kinematics (F), and visual gaze (V). The normalized model scores use the following color coding: dark green (0.980.980.980.981.001.001.001.00), light green (0.950.950.950.950.980.980.980.98), orange (0.900.900.900.900.950.950.950.95), and red ($<0.9$). Normalized model scores, defined in Section 2.3, are rounded to two decimal places, potentially resulting in two models having a score of $1.00$ (e.g., the other model’s score could be $0.997$).
@@
3.2 Fast- versus slow-timescale predictions are context-dependent
Biological motor control exhibits remarkable adaptability, seamlessly combining fast- and slow-timescale control strategies, depending on task complexity and other environmental demands. Here, we quantify multi-timescale tradeoffs as a function of varying terrains. We define the tradeoff between fast- and slow-timescale predictions as the relationship between CoM-relevant prediction intercept and its peak predictive power relative to the baseline. The intercept is defined by the $R^{2}$ value at gait phase 0 (Figure 3A), quantifying the degree of long-timescale prediction. We obtain the relative predictive power ($\Delta R^{2}$) by subtracting the baseline predictive power from the $R^{2}$, and define the peak relative predictive power as maximum $\Delta R^{2}$. While trials exhibit varying intercepts, the intercepts of CoM-relevant kinematics and the baseline are highly correlated (Figure 3B; $r=0.97$). By subtracting baseline predictive power from predictive power, this relative measure reduces inter-trial variability, enabling a more meaningful comparison across trials.
@@ 3.2 Fast- versus slow-timescale predictions are context-dependent
Figure 3: Relationship between modality-based intercept and peak relative predictive power. A Explained variance ($R^{2}$) as a function of gait phase for modality-based (orange) and baseline (blue) predictions. The intercept is defined as the modality-based $R^{2}$ value at gait phase 0 (i.e., the input time series spans from three gait cycles prior to touchdown and ends one gait cycle before touchdown). The red shaded area represents the relative predictive power ($\Delta R^{2}$), calculated as the difference between modality-based and baseline predictions. The peak is the maximum $\Delta R^{2}$. B Baseline intercept versus CoM-relevant kinematics intercepts for treadmill walking. C Relationship between the CoM-relevant kinematics intercepts and peak $\Delta R^{2}$ during overground walking across terrains with linear regression fits. The plots display the regression line and 95% confidence interval. D Coefficients between CoM-relevant kinematics $\Delta R^{2}$ peak and intercept of CoM-relevant kinematics. The box-plot shows the median (bar in box), $25-75\%$ percentile (box), and non-outlier range (whiskers) of the bootstrap statistics. The magnitude of the slope quantifies the strength of the tradeoff between long- and short-timescale prediction.
@@
3.3 Relative predictive power reveals different control timescales across modalities
@@
Human locomotion involves dynamic integration of multiple sensory inputs to guide future foot placement. Our framework seeks to uncover how humans utilize various input modalities, including CoM-relevant kinematics, full-body kinematics, and gaze fixations, across varying contexts like treadmill walking [7], treadmill running [37], and overground walking on terrains of different roughness [30]. To understand the predictive ability of the input modality, we train a model that takes the time series of the input modality and predicts the future foot placement. As the input time window increases, the model demonstrates improved predictive power, as shown by a closer alignment between predictions and ground truth (Figure 4A). The results demonstrate how different input modalities influence foot placement prediction across contexts. By analyzing the relative predictive power ($\Delta R^{2}$) of modality-based models compared to the baseline, we obtain the timing and magnitude of each modality’s contribution to foot placement control (Figure 4B).
@@
During overground walking on flat terrain, gaze does not play a significant role in foot placement prediction, as its performance does not surpass the baseline (Figure 4C). This suggests that foot placement control on flat terrain is not guided by gaze. Across contexts, full-body kinematics begin to outperform the baseline earlier than CoM-relevant kinematics; we provide some possible explanations in the discussion section. Despite the inclusion of significantly more body markers (29 vs. 1), full-body kinematics only explained 6-14% of additional foot placement variance, which highlights the significance of CoM-relevant kinematics in locomotor control (Figure 4C). As terrain complexity increases, foot placements are better predicted by gaze, with predictions earlier than both CoM-relevant and full-body kinematics for lateral (Figure 4C) but not fore-aft (Figure 8C) adjustments. This indicates that in more challenging environments, humans rely on visual gaze to inform lateral foot placement in advance of the body states. Across contexts and modalities, lateral foot placement control occurs earlier than fore-aft control, indicating a prioritization of lateral stability during locomotion (Figure 8A), which aligns with previous findings [7, 50, 51]. During treadmill running, the relative predictive power for lateral foot placement is not significantly different between even and uneven terrain. However, the fore-aft foot placement predictions is earlier during on uneven than on even terrain (Figure 8B). This discrepancy is likely a result of solely fore-aft terrain perturbations introduced in the uneven terrain running study [37].

### v3

@@ Abstract
Modeling movement in real-world tasks is a fundamental goal for motor control, biomechanics, and rehabilitation engineering. However, widely used data-driven models of essential tasks like locomotion make simplifying assumptions such as linear and fixed timescale mappings between past inputs and future actions, which do not generalize to real-world contexts. Here, we develop a deep learning-based framework for action prediction with architecture-dependent trial embeddings, outperforming traditional models across contexts (walking and running, treadmill and overground, varying terrains) and input modalities (multiple body states, gaze). We find that neural network architectures with flexible input history-dependence like GRU and Transformer perform best overall. By quantifying the model’s predictions relative to an autoregressive baseline, we identify context- and modality-dependent timescales. These analyses reveal that there is greater reliance on fast-timescale predictions in complex terrain, gaze predicts future foot placement before body states, and the full-body state predictions precede those by center-of-mass-relevant states. This deep learning framework for action prediction provides quantifiable insights into the control of real-world locomotion and can be extended to other actions, contexts, and populations.
keywords:
Machine Learning, Motion Prediction, Motor Control
Introduction
@@
Modeling human movement in real-world settings remains a major scientific challenge for motor control, biomechanics, and rehabilitation engineering. Indeed, data-driven models of essential tasks like locomotion have informed predictive simulations [1, 2], helped characterize neuromotor disorders [3, 4], and have been incorporated into wearable robots to anticipate the user’s actions [5, 6]. Data-driven models widely used to characterize locomotor control rely on simplifying assumptions such as linear, fixed-timescale mappings or are limited to laboratory-constrained tasks. Such models have successfully explained how center-of-mass (CoM) kinematics predict foot placement [7, 8, 9, 10] and ground reaction forces [2, 11] during treadmill locomotion, but they have not been shown to generalize to real world settings. During real world movements, people exhibit more variable behaviors [12, 13], and the control strategies may depend on the nonlinear history of multiple task-relevant inputs [14]. Addressing this requires a framework capable of predicting future actions given time-varying inputs from different modalities, such as body state and visual gaze, across contexts. Artificial neural networks have shown significant promise in predicting foot placement actions, as deep learning models can accurately capture the temporal and nonlinear statistics at play [15]. However, previous neural network models of locomotion test a single architecture [16, 17, 18], use a single input modality [16, 17] or are confined to a single laboratory-constrained setting like treadmill walking [18, 19]. Furthermore, these data-driven models do not separate the variance explained by the input modality from the variance explained by the inter-trial variability, hindering model performance and interpretability. Here, we put forth a deep learning framework comparing multiple neural network architectures, with architecture-dependent trial embeddings, outperforming existing models across real world contexts.
@@
Biological motor control is time-varying and context-dependent, requiring adaptive strategies to meet environmental, task-specific, and physiological demands [20]. For example, during locomotion, humans possess a remarkable ability to continuously modulate their actions in diverse “contexts" i.e. environment-specific circumstances such as uneven terrain or obstacle-avoidance [21, 22]. This ability is supported by dynamic sensorimotor control strategies that can be deployed both on a fast timescale such as during each walking step, or on a longer timescale over multiple steps [22, 23]. For instance, walking on uneven terrain requires rapid adjustments to foot placement [24], as well as persistent motor plans involving path or speed selection [25]. In fact, multiple timescales in motor control have been hypothesized to reflect distinct neural computations [26] or distinct functional goals [27], so quantifying multi-timescale control strategies from real world data has theoretical value. Despite this, we lack meaningful approaches to quantify multi-timescale control in real world contexts. Here, we put forth the predictive power of nonlinear action prediction models relative to an autoregressive baseline as a generalizable way to quantify control timescales, discovering a context-dependent tradeoff between fast- and slow-timescale predictions.
@@
Biological sensorimotor control is achieved by dynamically mapping multi-sensory inputs to motor commands, enabling adaptive control in changing contexts [28, 29]. During locomotion, the control of foot placement is inherently multimodal, and prior work has examined the role of individual input modalities such as vision [24] or low-dimensional body states on control [7]. Visual input is crucial for anticipatory locomotor control when navigating obstacles or selecting stable footholds on complex terrains [13, 30], while joint-level or CoM-based feedback informs muscle responses to transient perturbations [31, 32, 2]. Vestibular input is also thought to guide foot placement, particularly in a phase-dependent manner [33]. Despite the existence of multiple input modalities, existing foot placement control models assume a linear fixed control timescale with low-dimensional body states as an input [7, 34], thereby neglecting the temporal dynamics of how and when each modality contributes to control, and how different modalities compare to one another. To address these gaps, we propose a data-driven framework that systematically evaluates the contributions of multiple input modalities — including CoM-relevant kinematics, full-body kinematics, and gaze fixations — to predicting future foot placement. By comparing the predictive power of each modality to an autoregressive baseline model, we quantify the timescale and magnitude of modality-dependent foot placement. Our framework reveals at which phase of the gait cycle each input modality become critical for informing future foot placement.
@@
In this study, we develop deep learning-based models to predict future actions from multimodal input time series and test them in diverse locomotor tasks. By comparing different model architectures, we find that architectures with flexible input history-dependence generalize better across real-world contexts and input modalities. We characterize the control timescale of a given nonlinear action prediction model using its predictive power relative to an autoregressive baseline model, to help reveal how these timescales are shaped by the context and the input modality. Our work has implications for characterizing, modeling, and predicting real world movements and the multi-timescale control strategies hidden in them. This deep learning framework can inform the development of human-aware wearable robot control and more human-like physics-based simulations of motor behavior in the real world [35, 36].
Methods
@@
Data Description and Processing
@@
Figure 1: Overview of the data-driven modeling framework. A Inputs contain hypothesized modalities such as CoM-relevant kinematics, swing-foot kinematics, full-body kinematics, and gaze fixations (defined by the intersection of the gaze vector with the ground plane). The output is foot placement. B Foot placement output is defined by the lateral and fore-aft foot placement relative to the opposite foot at its previous heel strike. C The input is a time series spanning from three strides prior to touchdown to a specific gait phase within the previous stride. The output corresponds to foot placement at heel strike. We denote the ending gait phase of the time series relative to one stride prior to touchdown as $\phi$, discretized and sampled from $n=21$ equally spaced phases within $0\leq\phi\leq 1$. D The network architecture of the nonlinear models. The input consists of a time series, an L/R flag, and a trial ID. The time series and L/R flag are concatenated, while the trial ID is embedded to provide additional trial-specific information. The incorporation of this embedding is architecture-dependent, i.e. it influences the model prediction differently depending on the neural network architecture, as detailed in the Methods section.
Nonlinear models
@@
Model Evaluation and Baseline Model
@@
Nested cross-validation was employed to evaluate the models’ performance and hyperparameter tuning, ensuring an unbiased estimate of model performance and preventing data leakage [47]. The dataset was initially split into five equal folds. In each iteration of the outer loop, 80% of the data was allocated for training and validation, while the remaining 20% was held out as the testing set. The training/validation set was further subdivided into five inner folds. Four inner folds were used to train the model, and the remaining fold was used for early stopping, with a patience of 50 epochs. This process was repeated for every combination of hyperparameters, with parameter updates performed using the ADAM optimizer [48] and a learning rate of 0.001. We used mean squared error (MSE) as the loss function. After each training session, the model’s performance was evaluated on the outer testing fold. The outer loop was repeated five times, ensuring that each data point in the original dataset was used once for testing and four times for training/validation.
@@
Optimal hyperparameters for each modality and dataset were first identified using grid search (or random search for large hyperparameter spaces exceeding 100 configurations; see Appendix for details). These hyperparameters were then used in the nested cross-validation process. Aggregated predictions from all outer folds, covering 100% of the data points, were used to compute the explained variance ($R^{2}$) of foot placement predicted by the input modality. Under this nested cross-validation framework, it is possible, though unlikely, to obtain negative $R^{2}$ values, which occurs when the model fails to generalize to the testing dataset, performing worse than simply predicting the average foot placements (which would result in $R^{2}=0$). In such cases, the model’s predictive power is minimal or nonexistent. To prevent overfitting, we used early stopping and dropout when training the neural networks, whereas $L2$ regularization is used to prevent overfitting to historical data or high-dimensional input space for linear models. Since neural networks are more complex than linear models, they inherently exhibit lower bias and higher variance. To mitigate this higher variance of neural networks and to better interpret model behavior across gait phases, we smooth the $R^{2}$ curves using local regression (LOWESS) and cubic spline.
@@
To evaluate model performance, we assigned a score to each model based on the Root Mean Squared Error (RMSE). As the time window increases, the input contains accumulative information about the foot placement thereby enhancing the models’ predictive power. We used the RMSE curve achieved by the swing-foot kinematics as a baseline, reflecting the predictive information inherently contained in the swing foot dynamics [49]. We define the model score as the RMSE as defined in next section). The normalized model score is then computed relative to the model with the highest score. Formally, let $\text{RMSE}(m,\phi)$ denote the RMSE of model $m$ at gait phase $\phi$ and let $c$ denote the critical gait phase. We define the model score $s_{m}$:
@@
Quantifying the prediction timescale
The predictive power of foot placement using different input modalities across the gait cycle provides insight into the timescale of the control strategy. For an input modality to be considered a strong predictor, it must outperform the autoregressive baseline and demonstrate an earlier and larger increase in predictive power. To evaluate this, we identify the gait phase where the input modality achieves its maximum performance over the baseline, measured by the relative RMSE gap (Figure 2A). Beyond this phase, the input modality’s contribution diminishes, as the predictive power of the swing foot itself increases at a faster rate.
@@
Figure 2: Identification of key timepoints for dynamics and prediction. A The relative RMSE between modality-based predictions and the autoregressive baseline prediction as a function of gait phases. The timing of the peak difference indicates the phase where the input modality is the strongest predictor. B An example velocity curve along the locomotor axis, with the 5% of peak velocity threshold used to determine the onset of the swing phase. C The baseline $R^{2}$ curve, showing a breakpoint where $R^{2}$ increases significantly. The timing of this breakpoint is identified as the phase where the slope exceeds the average slope.
@@
We analyze the autoregressive baseline $R^{2}$ curve to determine the time at which the swing foot kinematics begins predicting substantially the future foot placement. The $R^{2}$ curve typically shows a non-decreasing trend as the gait cycle progresses, eventually reaching $R^{2}=1$ at the final timestep. A notable “breakpoint” in the $R^{2}$ curve is the phase at which the predictive power of the swing foot kinematics sharply rises. To systematically identify this breakpoint, we calculate the average slope of the $R^{2}$ curve, defined as $r=R^{2}(1)-R^{2}(0)$, and locate the gait phase where the slope reaches its maximal value. We then trace backward from this gait phase to find the first point where the slope falls below the average slope, marking the critical transition phase (Figure 2C). This approach provides a robust method for identifying key transitions in the gait cycle and aligns with the underlying dynamics of gait and motor control.
Results
@@
In this section, we present findings using our deep learning framework for action prediction to understand the influence of the environmental context and input modality on future foot placement. First, we demonstrate that neural network models, particularly GRUs, significantly outperform linear models in more challenging contexts. During overground walking, we find that subjects exhibit a tradeoff between fast and slow timescale predictions, with the magnitude of the tradeoff varying across different terrains. By comparing the predictive power of input modalities to an autoregressive baseline model, we reveal that control timescales are both modality- and context-dependent. Finally, we discover a gait phase mid-swing when the swing foot dynamics becomes useful for foot placement prediction.
@@
Models with flexible nonlinear input dependence outperform linear models
@@
We evaluated the performance of nine models — LSTM, GRU, FCNN, TCN, Transformer, Linear instance with (LI2) and without (LI) L2 regularization, Linear history with (LH2) and without (LH) L2 regularization — on their ability to predict foot placement across various tasks and input modalities. The tasks included treadmill walking, treadmill running, and overground walking on various terrains, while the input modalities consisted of CoM-relevant kinematics, full-body kinematics, swing-foot kinematics and gaze fixations. The model score was evaluated based on the Root Mean Squared Error (RMSE) relative to the optimal model at each gait phase, and then normalized by the highest-scoring model. The Methods section details the evaluation metric.
@@
TW N/A C 1.00 1.00 0.98 0.71 0.98 0.97 0.26 0.97 0.99
C 0.98 1.00 0.96 0.92 0.98 1.00 0.34 1.00 0.97
Even F 1.00 1.00 0.89 0.82 0.94 0.38 0.62 1.00 0.78
C 0.98 1.00 0.96 0.91 0.97 0.97 0.31 0.97 0.97
TR Uneven F 1.00 1.00 0.93 0.91 0.96 0.35 0.63 0.99 0.76
C 0.96 1.00 0.95 0.94 0.99 0.91 0.70 0.95 0.88
F 0.97 0.99 0.94 0.95 1.00 0.52 0.76 0.87 0.77
Flat V 0.80 0.87 0.70 0.91 1.00 0.69 0.30 0.75 0.30
C 0.96 0.99 0.95 0.98 1.00 0.86 0.66 0.93 0.78
F 0.97 0.98 0.94 0.99 1.00 0.52 0.76 0.82 0.77
Medium V 0.95 0.98 0.88 1.00 0.97 0.68 0.32 0.75 0.33
C 0.99 1.00 0.97 0.99 1.00 0.31 0.68 0.87 0.73
F 0.99 0.99 0.94 0.98 1.00 0.58 0.74 0.74 0.74
OW Rough V 1.00 1.00 0.92 0.96 0.97 0.34 0.61 0.61 0.61
Table 1: Normalized Model Scores across Tasks and Input Modalities. Models were evaluated on treadmill walking (TW), treadmill running (TR), and overground walking (OW) tasks with varying terrain. The input modalities (M) were: CoM-relevant kinematics (C), full-body kinematics (F), and visual gaze (V). The normalized model scores use the following color coding: dark green (0.980.980.980.981.001.001.001.00), light green (0.950.950.950.950.980.980.980.98), orange (0.900.900.900.900.950.950.950.95), and red ($<0.9$). Normalized model scores, defined in the Methods section, are rounded to two decimal places, potentially resulting in two models having a score of $1.00$ (e.g., the other model’s score could be $0.997$).
@@
Fast- versus slow-timescale predictions are context-dependent
Biological motor control exhibits remarkable adaptability, seamlessly combining fast- and slow-timescale control strategies, depending on task complexity and other environmental demands. Here, we quantify multi-timescale tradeoffs as a function of varying terrains. We define the tradeoff between fast- and slow-timescale predictions as the relationship between CoM-relevant prediction intercept and its peak predictive power relative to the autoregressive baseline. The intercept is defined by the $R^{2}$ value at gait phase 0 (Figure 3A), quantifying the degree of long-timescale prediction. We obtain the relative predictive power ($\Delta R^{2}$) by subtracting the baseline predictive power from the $R^{2}$, and define the peak relative predictive power as maximum $\Delta R^{2}$. While trials exhibit varying intercepts, the intercepts of CoM-relevant kinematics and the baseline are highly correlated (Figure 3B; $r=0.97$). By subtracting baseline predictive power from predictive power, this relative measure reduces inter-trial variability, enabling a more meaningful comparison across trials.
@@
Figure 3: Relationship between modality-based intercept and peak relative predictive power. A Explained variance ($R^{2}$) as a function of gait phase for modality-based (orange) and autoregressive baseline (blue) predictions. The intercept is defined as the modality-based $R^{2}$ value at gait phase 0 (i.e., the input time series spans from three gait cycles prior to touchdown and ends one gait cycle before touchdown). The red shaded area represents the relative predictive power ($\Delta R^{2}$), calculated as the difference between modality-based and baseline predictions. The peak is the maximum $\Delta R^{2}$. B Baseline intercept versus CoM-relevant kinematics intercepts for treadmill walking. C Relationship between the CoM-relevant kinematics intercepts and peak $\Delta R^{2}$ during overground walking across terrains with linear regression fits. The plots display the regression line and 95% confidence interval. D Coefficients between CoM-relevant kinematics $\Delta R^{2}$ peak and intercept of CoM-relevant kinematics. The box-plot shows the median (bar in box), $25-75\%$ percentile (box), and non-outlier range (whiskers) of the bootstrap statistics. The magnitude of the slope quantifies the strength of the tradeoff between long- and short-timescale prediction.
@@
Relative predictive power reveals different control timescales across modalities
@@
Human locomotion involves dynamic integration of multiple sensory inputs to guide future foot placement. Our framework seeks to uncover how humans utilize various input modalities, including CoM-relevant kinematics, full-body kinematics, and gaze fixations, across varying contexts like treadmill walking [7], treadmill running [37], and overground walking on terrains of different roughness [30]. To understand the predictive ability of the input modality, we train a model that takes the time series of the input modality and predicts the future foot placement. As the input time window increases, the model demonstrates improved predictive power, as shown by a closer alignment between predictions and ground truth (Figure 4A). The results demonstrate how different input modalities influence foot placement prediction across contexts. By analyzing the relative predictive power ($\Delta R^{2}$) of modality-based models compared to the autoregressive baseline, we obtain the timing and magnitude of each modality’s contribution to foot placement control (Figure 4B).
@@
During overground walking on flat terrain, gaze does not play a significant role in foot placement prediction, as its performance does not surpass the autoregressive baseline (Figure 4C). This suggests that foot placement control on flat terrain is not guided by gaze. Across contexts, full-body kinematics begin to outperform the baseline earlier than CoM-relevant kinematics; we provide some possible explanations in the discussion section. Despite the inclusion of significantly more body markers (29 vs. 1), full-body kinematics only explained 6-14% of additional foot placement variance, which highlights the significance of CoM-relevant kinematics in locomotor control (Figure 4C). As terrain complexity increases, foot placements are better predicted by gaze, with predictions earlier than both CoM-relevant and full-body kinematics for lateral (Figure 4C) but not fore-aft (Figure 8C) adjustments. This indicates that in more challenging environments, humans rely on visual gaze to inform lateral foot placement in advance of the body states. Across contexts and modalities, lateral foot placement control occurs earlier than fore-aft control, indicating a prioritization of lateral stability during locomotion (Figure 8A), which aligns with previous findings [7, 50, 51]. During treadmill running, the relative predictive power for lateral foot placement is not significantly different between even and uneven terrain. However, the fore-aft foot placement predictions is earlier during on uneven than on even terrain (Figure 8B). This discrepancy is likely a result of solely fore-aft terrain perturbations introduced in the uneven terrain running study [37].